OpenStudy (anonymous):

the point p(x,y) on the unit circle that corresponds to a real number t is given. find the value of the indicated trigonometric function at t.

OpenStudy (anonymous):

\[(-\sqrt{65}/9 , -4/9)\]

OpenStudy (anonymous):

find sin t

OpenStudy (anonymous):

Correct me if I'm wrong but is this using t-results or something?

OpenStudy (anonymous):

(-sqr65/9, -4/9) find sin t

OpenStudy (anonymous):

You should be drawing these questions out so you can graphically see what's going on.

jimthompson5910 (jim_thompson5910):

sin(t) = -4/9 arcsin(sin(t)) = arcsin(-4/9) t = arcsin(-4/9) t = ???